### v1

# ChainRank-DPO: Chain Rank Direct Preference Optimization for LLM Rankers

Abstract
Large language models (LLMs) have demonstrated remarkable effectiveness in text reranking through works like RankGPT, leveraging their human-like reasoning about relevance. However, supervised fine-tuning for ranking often diminishes these models’ general-purpose capabilities, including the crucial reasoning abilities that make them valuable for ranking. We introduce a novel approach integrating Chain-of-Thought prompting with an SFT-DPO (Supervised Fine-Tuning followed by Direct Preference Optimization) pipeline to preserve these capabilities while improving ranking performance. Our experiments on TREC 2019 and 2020 Deep Learning datasets show that our approach outperforms the state-of-the-art RankZephyr while maintaining strong performance on the Massive Multitask Language Understanding (MMLU) benchmark, demonstrating effective preservation of general-purpose capabilities through thoughtful fine-tuning strategies. Our code and data will be publicly released upon the acceptance of the paper.
ChainRank-DPO: Chain Rank Direct Preference Optimization for
LLM Rankers
Haowei Liu Santa Clara University Santa Clara, CA hliu6@scu.edu                        Xuyang Wu Santa Clara University Santa Clara, CA xwu5@scu.edu                        Guohao Sun Rochester Institute of Technology Rochester, NY gs4288@rit.edu
Zhiqiang Tao Rochester Institute of Technology Rochester, NY zhiqiang.tao@rit.edu                        Yi Fang Santa Clara University Santa Clara, CA yfang@scu.edu
1 Introduction
Text reranking is a vital task in information retrieval Liu (2009); Hasanain (2018), crucial for search engines Li et al. (2024), conversational AI Becker et al. (2012), and recommendation systems Zhao and Liu (2024). Large language models (LLMs) excel in reranking due to their reasoning and human-like thinking capabilities, enabling them to handle complex queries and ambiguous contexts. RankGPT Achiam et al. (2023) has set a high standard in listwise reranking Ma et al. (2019), leveraging its interpretability and systematic reasoning Brown et al. (2020) to advance reranking technologies.
Figure 1: While both methods can rerank passages, general methods struggle with simple QA tasks, whereas ChainRank successfully solves them step by step.
Building on the success of RankGPT, several models have sought to distill its output into smaller, task-optimized models through supervised fine-tuning (SFT). RankVicuna Pradeep et al. (2023a), for instance, uses distilled data generated exclusively from GPT-3.5, focusing on efficient fine-tuning to improve ranking performance while maintaining a lightweight architecture. Moreover, RankZephyr Pradeep et al. (2023b) employs a more comprehensive approach utilizing distilled data from GPT-3.5 and GPT-4, ensuring the quality and reliability of training data to achieve superior ranking results. Although these models achieve state-of-the-art performance in specific benchmarks, they highlight a key challenge in ranking: the trade-off between optimizing task-specific performance and preserving the broader reasoning capabilities of LLMs. Addressing this challenge requires novel strategies to balance specialization with generalization, ensuring that ranking models retain their versatility across diverse tasks.
Our investigations in Figure 1 reveal that while RankVicuna and RankZephyr exhibit strong performance in listwise ranking tasks, they have forfeited their mathematical reasoning capabilities following fine-tuning. This is evident from the figure, where the models fail to produce coherent solutions to mathematical problems, often outputting unrelated or nonsensical numerical sequences. Moreover, further analysis uncovered that these models have also lost their general text-generating capabilities. Regardless of the input prompt, the models struggle to produce meaningful or contextually appropriate responses, highlighting a significant trade-off introduced by their specialized fine-tuning for ranking tasks. Furthermore, while they exhibit comprehension of listwise ranking, they encountered difficulties with pairwise ranking, as their pairwise outcomes did not correspond with those generated by the listwise ranking Peng et al. (2024).
To address this limitation, we propose ChainRank strategy, built on LLaMA3-8b-instruct Dubey et al. (2024). Using the same training data as RankZephyr and RankGPT3.5/RankGPT4 as teacher models, we introduced a chain-of-thought (CoT) prompt to guide sequential passage ranking. We then implemented ChainRank-DPO, enhancing the model’s reasoning abilities for superior ranking performance. While traditional Direct Preference Optimization (DPO) Rafailov et al. (2024) performs well on chat benchmarks, it struggles with long-chain reasoning tasks like math and ranking due to error propagation. Inspired by Step-DPO Lai et al. (2024), we designed a novel DPO framework using overlapping ranking orders as the reward function, allowing better error correction. ChainRank preserves general text generation and reasoning capabilities, overcoming limitations in RankVicuna and RankZephyr. It outperforms both on TREC 2019/2020 Deep Learning Tracks Craswell et al. (2021, 2020) and BEIR benchmarks Thakur et al. (2021). We summarize the contributions of this work as follows:
  • We propose a novel chain-of-thoughts instruction (reranking) tuning that enables LLMs to rank passages based on relevance step by step.
  • To our best knowledge, this is the first work to introduce the Supervised Fine-Tune–DPO pipeline in the context of text reranking.
  • Using the same training data, our model outperforms both RankZephyr and the teacher model, RankGPT4, while retaining the general capabilities of LLMs.
{mdframed}
[ linecolor=black!60, linewidth=1pt, roundcorner=10pt, backgroundcolor=gray!5, shadow=true, shadowsize=5pt, shadowcolor=black!40, skipabove=10pt, skipbelow=10pt, innertopmargin=10pt, innerbottommargin=10pt, innerleftmargin=10pt, innerrightmargin=10pt ] USER: I will provide you with {num} passages, each indicated by a numerical identifier []. Rank the passages based on their relevance to the search query: {query}.
[1] {passage 1}
[2] {passage 2}
[{num}] {passage {num}}
Search Query: {query}.
Rank the {num} passages by selecting the most relevant passage at each step from the remaining passages. After choosing the most relevant passage, remove it from the pool and continue ranking until all passages are ordered.
Instructions:
Start with the most relevant passage and select it from the full list.
For each following step, pick the most relevant passage from the remaining passages only.
List the selected passages by their identifiers at each step, one after the other, until all passages are ranked.
Example Output:
Step 1: [4]
Step 2: [4, 2]
Step 3: [4, 2, 3]
step {num}: [4,2,3,15,…,14]
Final Answer: [4, 2, 3,…, 14]
Only respond with each step and the final answer, ensuring each passage is included once and ranked in descending relevance.
Figure 2: ChainRank Chain-of-Thought (CoT) reranking prompt guiding the model to rank passages based on relevance to a query iteratively. The prompt ensures step-by-step selection, removal, and ordering of passages, with an example illustrating the expected output format.
2 Methodology
Figure 3: Training framework of our model. $Left:$ Stage 1 aims to teach the student model to perform text reranking. $Right:$ In stage 2, we use the same prompt format to generate multiple answers, then pick the chosen and rejected answers based on the number of overlapped steps, finally, we utilize the preference data to perform Chain DPO training.
2.1 CoT Reranking Prompt Design
Recent advancements in listwise text reranking, such as RankVicuna Pradeep et al. (2023a), RankZephyr Pradeep et al. (2023b), and related methods Liu et al. (2024), build on zero-shot techniques Ma et al. (2023); Sun et al. (2023), using prompt templates to reorder documents for metrics like nDCG Järvelin and Kekäläinen (2002). Given a query $q$ and documents ${d_{1},d_{2},...,d_{n}}$, the goal is to return a reordered list (e.g., $[1]>[2]>[3]...>[20]$).
Our ChainRank strategy frames listwise reranking as a chain-of-thought (CoT) reasoning task, selecting the most relevant document iteratively until all are ranked. The CoT prompt for LLaMA is shown in Figure 2.
2.2 Training Dataset
We train ChainRank model using 35k GPT-3.5 and 5k GPT-4 labeled instances from Pradeep et al. (2023b), derived from randomly selected MS MARCO v1 queries. Pyserini Lin et al. (2021) retrieved 20 BM25 candidates per query, which RankGPT3.5 and RankGPT4 ordered as teacher models. Malformed generations, such as missing or duplicate identifiers, were removed to improve data quality. We randomly selected 90% of the query-document pairs for Stage 1 SFT training and 10% for Stage 2 DPO training, denoted as SFT data and DPO data.
2.3 Training Stage 1: Supervised Fine-Tune
As shown in Figure 3, during the Supervised Fine-Tuning (SFT) phase, we use SFT data, maintaining a zero-shot setup since RankGPT and ChainRank do not rely on human-labeled data.
Before fine-tuning, the original LLaMA3-8B-Instruct model Dubey et al. (2024) failed to produce meaningful CoT reranking, simply replicating the example format without relevance-based results, highlighting the need for fine-tuning. The model weights are publicly available on HuggingFace. [^1] https://huggingface.co/meta-LLaMA/Meta-LLaMA-3-8B-Instruct
We fully fine-tune the 8B parameter LLaMA-3 model for three epochs with a batch size of 128, a $5\times 10^{-6}$ learning rate, and bfloat16 format. Training on four NVIDIA A100 80GB GPUs took approximately 39 hours.
2.4 Training Stage 2: Chain DPO
After Stage 1, the ChainRank-SFT model generates three ranking predictions ($y$) on prompts ($x$) from DPO training data. Predictions are evaluated by overlapping ranking orders with ground-truth labels, creating a preference dataset with prompts (x), chosen steps ($s_{w}$), rejected steps ($s_{l}$), and overlapping steps ($s_{o}$). Unlike prior methods, the final ranking $y$ comprises a sequence of reasoning steps, $y={s_{1},s_{2},\dots,s_{n}}$, where each step is conditioned on prior steps $\pi(s_{k}|x;s_{1:k-1})$. Overlapping steps ($s_{o}$) are tracked until a divergence is found, after which only the initial contiguous overlaps are included in $s_{o}$, and later steps are categorized as chosen ($s_{w}$) or rejected ($s_{l}$). This results in a dataset of $(x,s_{w},s_{l},s_{o})$.
The objective maximizes the likelihood of correct steps ($s_{w}$) while minimizing incorrect ones ($s_{l}$), using the loss:
$\begin{split}&\mathcal{L}(\theta)=-\mathbb{E}_{(x,s_{w},s_{l},s_{o})\sim D}\Big{[}\log\sigma\Big{(}\\ &\beta\log\frac{\pi_{\theta}(s_{w}\mid x;s_{o})}{\pi_{ref}(s_{w}\mid x;s_{o})}-\beta\log\frac{\pi_{\theta}(s_{l}\mid x;s_{o})}{\pi_{ref}(s_{l}\mid x;s_{o})}\Big{)}\Big{]}\,,\end{split}$
where $\pi_{\theta}(s\mid x;s_{o})$ represents the model’s probabilities, optimized to favor correct steps while leveraging reference model probabilities ($\pi_{\text{ref}}$) for stable training and better generalization, especially in zero-shot scenarios. This improves CoT reasoning’s sequential decision-making and model consistency. The Chain DPO stage is trained on four NVIDIA A100 80GB GPUs over one epoch, taking approximately eight hours.
3 Experiments
3.1 Research Questions
Before conducting the experiments, we formulate several research questions to ensure that ChainRank could effectively address the challenges:
-RQ1: Does the CoT instruction tuning improve the step-by-step relevance ranking of passages compared to traditional ranking methods?
-RQ2: How does our model’s performance on text reranking tasks compare with existing models like RankZephyr and RankGPT4 across various datasets and task settings?
-RQ3: Does introducing DPO improve the performance and robustness of our model?
3.2 Evaluation Benchmarks
We evaluate ranking capabilities using TREC DL19 Craswell et al. (2020) and DL20 Craswell et al. (2021) Tracks, derived from MS MARCO V1 Bajaj et al. (2016), with human-annotated relevance labels. TREC DL19 and DL20 contain 43 and 54 queries, respectively, each paired with 100 candidate passages retrieved by BM25. Additionally, we include cross-domain datasets from BEIR Thakur et al. (2021), not seen by the fine-tuned LLMs, for broader analysis. Using a sliding window strategy with a size of 20 and a stride of 10, we rerank 100 passages per query.
We use the Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2020) to evaluate whether our model retains its text understanding abilities after fine-tuning for ranking tasks. MMLU tests proficiency across 57 diverse subjects, from elementary to advanced professional topics. We compare our model’s performance on MMLU with baseline LLMs to ensure ranking fine-tuning does not compromise general-purpose language understanding and generation capabilities, maintaining its versatility in broader tasks.
3.3 Results
Table 1 compares model performance on TREC, BEIR, and MMLU benchmarks. Our 8B ChainRank model outperforms baselines, including RankGPT4  [^2] RankGPT is the only closed-source and largest model in the table., even with 90% of the training data, and surpasses RankLLaMA3 under identical training conditions, showcasing the effectiveness of our CoT reranking prompt. In the DPO stage, ChainRank-DPO improves across all datasets with just one training epoch, enhancing robustness to passage variations (Appendix B).
On MMLU, ChainRank matches LLaMA3’s performance, while RankLLaMA3 shows a slight drop and RankVicuna performs poorly. RankZephyr has completely lost its ability to generate meaningful outputs, it receives a score of 0 on MMLU. Additional examples are in Appendix D.
We present a figure to illustrate the trade-off between performance and inference cost when generating ranking orders at varying step intervals. Notably, our method achieves comparable latency to RankLLaMA3 (26.5 vs. 25.0 seconds). By utilizing parallel distributed evaluation, we further reduce inference latency to 6.73 seconds on 4 A100 GPUs, demonstrating the efficiency of our approach.
Models TREC BEIR MMLU DL19 DL20 NFC COVID FIQA AVG. RankGPT4 0.746 0.708 0.406 0.749 0.333 0.864 BM25 0.506 0.480 0.325 0.595 0.236 N/A Contriever 0.616 0.599 0.328 0.596 0.329 N/A Gemma-7B 0.533 0.530 0.338 0.573 0.236 0.649 LLaMA3 0.641 0.621 0.269 0.628 0.216 0.662 RankZephyr 0.742 0.709 0.331 0.592 0.230 0.000 RankVicuna 0.668 0.655 0.338 0.592 0.236 0.373 RankLLaMA3 0.730 0.635 0.335 0.672 0.330 0.628 ChainRank-SFT 0.752 0.714 0.353 0.766 0.342 0.663 ChainRank-DPO 0.755 0.717 0.358 0.772 0.342 0.663
Table 1: Performance of different models on TREC (nDCG@10) and MMLU (exact match score) benchmarks. All the reranking tasks are based on BM25 retrieval results. The bold values in the table highlight the best performance across the respective benchmarks.
Figure 4: FLOPs and nDCG@10 performance across different step intervals. The blue line represents the inference cost (FLOPs), while the orange line shows the ranking performance (NDCG@10). As the step interval increases, FLOPs decrease significantly, with a slight drop in nDCG@10 scores.
4 Conclusion and Future Work
In this paper, we propose a novel approach, ChainRank, a novel zero-shot listwise text reranking model built on LLaMA3. We demonstrated the effectiveness of CoT reranking prompt and our SFT–Chain DPO pipeline. The results of our experiments show that our model achieves superior performance compared to many open-source and closed-source LLMs, such as RankZephyr and RankGPT4.
In future work, we plan to explore the application of our method to other models, such as Mistral, Zephyr, and LLaMA3.1, to evaluate its generalizability and performance across different architectures. Additionally, our future research could incorporate higher-quality datasets with varying numbers of passages per instance to enhance the diversity and robustness of the training data. Expanding the dataset in this manner could lead to improved overall performance.
5 Limitations
While this study provides valuable insights, several limitations should be acknowledged. Firstly, the need to include long examples of CoT formatting (Step1: …) in our prompts leads to increased training and inference times.
Secondly, although LLaMA3 has a larger context window (8192 tokens) compared to Vicuna and Zephyr (4096 tokens), we are still limited by the maximum number of passages in our training data, which is set to 20. As a result, we are restricted to ranking within a window size of 20 and require nine passes to complete the reranking of 100 passages.
Appendix A Baseline Selection
To illustrate the efficacy of ChainRank, we select several baseline models, including the existing retriever models, BM25 Robertson and Zaragoza (2009) and Contriever Izacard et al. (2021), zero-shot prompt-based LLMs, RankGPT4 and Gemma-7B Team et al. (2024), fine-tuned LLM-based models (RankVicuna, RankZephyr), as well as our base model LLaMA3-8B-Instruct. To show the effectiveness of our CoT reranking prompt, we train a model with a similar paradigm to RankVicuna but use our base model LLaMA3-8B-Instruct, denoted as RankLLaMA3.
Appendix B Window Size
Models Window/Stride Size 2/1 10/5 20/10 Gemma-7B 0.541 0.593 0.533 LLaMA3 0.549 0.660 0.641 RankZephyr 0.562 0.612 0.742 RankVicuna 0.551 0.655 0.668 RankLLaMA3 0.567 0.663 0.666 ChainRank-SFT 0.697 0.737 0.755 ChainRank-DPO 0.671 0.755 0.758
Table 2: Performance of different models with different window size and step size, BM25 performs all of the retrieval stage, and all of the evaluations are performed on DL19
Appendix C Explanation of CoT strategies
To evaluate the trade-off between performance and inference cost, we trained our model to generate ranking orders after every 1, 3, 5, and 7 steps. An additional instruction was introduced to guide the model in returning ranking orders at these specified intervals. For comparison, we also included the performance of RankLLaMA3 to highlight the effectiveness of our CoT strategies
Appendix D Examples of general LLM tasks
To illustrate the degradation in general text understanding and generation capabilities of RankVicuna and Rankzephyr, we provide two examples below. Due to the length of ChainRank’s response, we have omitted the middle portion of the answer.
D.1 Example 1
Prompt Compose an engaging travel blog post about a recent trip to Hawaii, highlighting cultural experiences and must-see attractions.
RankVicuna Ultimately, showcasing Hawaii’s rich history and cultural significance. Title: Embracing Hawaiian Culture: Discovering Historical Landmarks and Rich Traditions on the Big Island
RankZephyr [1] > [2] > [3] > [4] > [5] > [6] > [7] > [8] > [9] > [10] > [2] > [11] > [12] > [13]
ChainRank Aloha and Mahalo: Discovering the Aloha Spirit in Hawaii As I stepped off the plane in Honolulu, I couldn’t help but feel a sense of excitement and anticipation wash over me. I had dreamed of visiting Hawaii for years, and finally, I was here. The warm tropical air enveloped me, and the sweet scent of plumeria filled my senses as I made my way to my hotel. Little did I know, this trip would be a journey not only to breathtaking natural wonders but also to a rich cultural heritage that would leave me feeling grateful and inspired. Immersing in the Local Culture: (Content omitted for brevity.) Must-See Attractions: (Content omitted for brevity.) A Taste of Aloha: (Content omitted for brevity.) Conclusion: (Content omitted for brevity.) Practical Tips and Recommendations: Visit Pearl Harbor and USS Arizona Memorial in the morning to avoid the crowds. Try local specialties like poke, laulau, and shave ice. Take a surf lesson or snorkeling trip to explore the islands’ stunning coastlines. Visit the Bishop Museum to learn about the history and geology of Hawaii. Explore the islands’ many farmers’ markets for fresh produce and local products. Learn traditional Hawaiian crafts, like leis and ukulele strings, from local artisans. Mahalo (thank you) for joining me on this journey to the Aloha State. I hope my story has inspired you to plan your own trip to Hawaii and experience the warmth and beauty of this incredible destination.
D.2 Example 2
Prompt Can you help me write a resignation letter to my current employer, while leaving on good terms and expressing gratitude for the opportunities provided?
RankVicuna (Nothing was generated from the model.)
RankZephyr [1] > [2] > [3] > [4] > [5] > [6] > [7] > [8] > [9] > [10] > [2] > [3] > [4] > [5] > [6]
ChainRank I’d be happy to help you write a resignation letter to your current employer. Here’s a sample template you can use as a guide: Dear [Recipient’s Name], I am writing to inform you of my decision to resign from my position as [Your Position] at [Company Name], effective [Date of Last Day of Work]. I want to express my deepest gratitude to you and the entire [Company Name] team for the opportunities I’ve had during my time here. (Content omitted for brevity.) Please don’t hesitate to reach out to me if you need anything after my departure. I can be reached at [Your Email Address] or [Your Phone Number]. Sincerely, [Your Signature] Some tips to keep in mind when writing your resignation letter: Be professional and respectful: Use a formal tone and language throughout the letter. Express gratitude: Let your employer know how much you appreciate the opportunities you’ve had and the experience you’ve gained. Offer to help with the transition: Show that you’re committed to making the transition as smooth as possible by offering to complete outstanding tasks and assist with the handover of your responsibilities. Keep it brief: Aim for a letter that is no more than one page in length. Proofread: Make sure to proofread your letter carefully to ensure it is free of errors and typos. Remember to customize the letter to fit your own circumstances and experiences. Good luck with your resignation, and I hope it leads to a positive and successful transition for you!